the ionization potential of H like species is 54.4ev the energy required to remove an electron from the n=2 state in this ion is

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Energy of electron in nth shell of a H like species of atomic number Z is

E\propto \dfrac{Z^2}{n^2}
ThereforeE_n=-54.4eV\dfrac{1^2}{2^2}=-13.6eV
So the required energy to remove the particular electron is 13.6eV.